Session: Jun 2006 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click

Experiment Feedback

Shanghai (TaoAn at Thu Jun 15 16:08:17 2006) Success :

Urumqi (wenjun yang at Thu Jun 15 16:21:27 2006) Minor failures : HIGH FS
It is rainy almost whole day, and meanwhile, the fs computer was dead, lost some data about 20min, more details is in log file. other is no problem.

Onsala (Michael Lindqvist at Thu Jun 15 17:57:11 2006) Success :
No known problems.

Torun (K.Borkowski at Fri Jun 16 05:20:42 2006) Success :
No known problems.

Westerbork (Tony Foley at Fri Jun 16 11:26:39 2006) Minor failures :
Wb used in phased array mode. Some formatter warnings (code 0x1004) at 04:47, 09:23 and 15:29UT. I think that this is just the UART busy. Also scan_check for scans 56,157,229, 293, 333 and 341 giving bytes lost and scans 181 and 261 giving "?,?,?" in the listing. Other scan_checks OK. TP jumps about 09:40UT in BBC2 (and both IFDs)

Noto (P. Cassaro at Sat Jun 17 12:17:34 2006) Success :
No known problems

Effelsberg (A. Polatidis at Mon Jun 19 08:45:06 2006) Minor failures : WIND
Observations stopped between 0500-0700 due to lightnings

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:25:23 2006) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Jun 20 09:26:07 2006) Minor failures : RECO
Data throttling reported on disk packs NTO-0017, WSRT-040 and WSRT-0046, although 1Gbps tests were passed. No replacements were available. Lost about 3 minutes from 15:00 UT due to restart of Mk5. No other known problems.

Medicina (Andrea Orlati at Tue Jun 20 09:27:14 2006) Minor failures : RECO ANT
Communication between ACU and Station computer fell down. These caused the FS to freeze. Lost scans between 06:30 UT and 7:40 UT even if the recording was not turned off(scan #179 included). 16:00 UT a bank_switch failed (scan between 14 and 14:21 lost). The only possible cause we saw is that the full module had a bad vsn label (there were strange characters). This also caused that the label stored into the log file after the bank switch (10:20 UT) was not rigth: jive-064 is reported but wsrt-037 is the correct one.

EVN experiment feedback facility: entire session

Entire Session

Session: Jun 2006 (session 2)
Wavelengths: 30cm 18/21 5 6 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Jun 5 12:32:29 2006) Minor failures :
This note concerns the UHF experiments only. Shortly before the first experiment we could manage to install the Lange coupler to produce two circular polarizations instead of the linear polarization announced in the block schedule. Also, we failed to install the phase cal.

Jodrell_Bank (Alastair Gunn at Tue Jun 20 09:14:12 2006) Success : PCAL
The MkII telescope was used (Jb2). During this session the Mk2 phase cal. signal was found to be too high and could not be easily reduced in strength. I therefore switched off the phas cal. signal for the entire session (for Mk2 only).